one number is 5 more than another. if the sum of twice the smaller number and three times the larger number is 75, find the two number
Answer by stanbon(75887)

one number is 5 more than another. if the sum of twice the smaller number and three times the larger number is 75, find the two number
------
1st #: x
2nd #: x+5
----
Equation:
2x + 3(x+5) = 75
2x + 3x + 15 = 75
5x = 60
x = 12
x + 5 = 17
============================
Cheers,
Stan H.
===========================